OpenStudy (anonymous):

Factor the polynomial. x3 + 2x2 + 4x + 8 Hint: use " i " to enter factors containing square roots of negative numbers. For example, if your factor is (x + ), you should write (x + 2i

OpenStudy (anonymous):

$$x^3+2x^2+4x+8=x^2(x+2)+4(x+2)=(x+2)(x^2+4)=(x+2)(x+2i)(x-2i) $$

OpenStudy (anonymous):

Thanxx Katarina .. Btw Simbasocks Mathway doesnt work with this problem

OpenStudy (anonymous):

U can see by eye that -2 is a root (use Rational Roots Theorem if u can't see it, gives a hit). Then solve the quadratic.

OpenStudy (anonymous):

(x+2)( ) Come on, not hard...

OpenStudy (anonymous):

okay thanxx but no need for the rudeness :)

OpenStudy (anonymous):

I thought I was being helpful:-( think I will stay clear of u in future.